### v2

# Electricity Market Bidding for Renewable Electrolyzer Plants: An Opportunity Cost Approach

Abstract
Hydrogen produced through electrolysis with renewable power is considered key to decarbonize several hard-to-electrify sectors. This work proposes a novel approach to model the active electricity market participation of co-located renewable energy and electrolyzer plants, based on opportunity-cost bidding. While a renewable energy plant typically has zero marginal cost, selling power to the grid carries a potential opportunity-cost of not producing hydrogen when it is co-located with a hydrogen electrolyzer. We first consider only the electrolyzer, and derive its revenue of consuming electricity based on the non-convex hydrogen production curve. We then consider the available renewable energy production and form a piece-wise linear cost curve representing the opportunity cost of selling (or revenue from consuming) various levels of electricity. This cost curve can be used to model a stand-alone electrolyzer or a co-located hydrogen and renewable energy plant participating in an electricity market. Our case study analyzes the effects of market-bidding electrolyzers on electricity markets and grid operations. We compare two strategies for a co-located electrolyzer-wind plant; one based on the proposed bid curve and one with a more conventional fixed electrolyzer consumption. The results show that electrolyzers that actively participate in the electricity market lower the average cost of electricity and the amount of curtailed renewable energy in the system compared with a fixed consumption case. However, the difference in total system emissions between the two strategies is insignificant. The specific impacts vary based on electrolyzer capacity and hydrogen price, which determines the location of the co-located plant in the electricity market merit order.
I Introduction
The deteriorating effects of climate change demand decarbonization efforts in all sectors. Certain sectors, such as chemical manufacturing, industrial heat and heavy transportation, are difficult to directly electrify. However, these sectors can be decarbonized through electrolytic hydrogen production powered by renewable energy sources (RES).
Electrolytic hydrogen is produced through “water-splitting”, where an electric current is applied to an electrolyzer, producing hydrogen (and oxygen) gas from water. The emissions of this process are mainly associated with the consumed power, and are close to zero if renewable power is consumed. The hydrogen can be further processed into a range of chemicals such as ammonia or methanol, which can be used, e.g., for the production of fertilizers or renewable fuels. The world already consumes around 95 Mt of hydrogen gas per year, although this hydrogen is mainly produced through reforming of natural gas, with emission intensities of around 10 kg CO2 eq. per kg H2 [1]. To decarbonize this hydrogen, governments are planning for the integration of large-scale electrolyzers into the power system, as laid out in the U.S. and E.U. strategies for hydrogen production [2, 3].
Although hydrogen produced through electrolysis enables low-carbon fuels and chemicals, the build-out of electrolyzers will cause significant increases in power demand [4]. Theoretically, electrolyzers can be flexible and controllable loads, and could fit well in modern power systems with high penetrations RES and limited transmission capacities [5]. In particular, if electrolyzers align their electricity consumption with the availability of renewable energy, their power demand might contribute to lift power prices and reduce the curtailment of renewable power, supporting the integration of renewable generation and improving the business case for continued expansion of renewable generation resources. However, while electrolyzers could theoretically be flexible, it is unclear if existing technology and economic incentives can encourage them to use this flexibility in a way that supports renewable generation. To mitigate these potential downsides, both the E.U. and U.S. have incentives that seek to tie hydrogen production to renewable generation [6, 7]. The primary goal of electrolyzers is to maximize profits from hydrogen production, creating a strong incentive to operate at high utilization rates to amortize the high capital costs associated with their construction. Further, electrolyzers might operate with Power Purchase Agreements (PPAs) and utility contracts that reduce their sensitivity to the real-time electricity price. Electrolyzers might therefore act more like large, constant loads than flexible resources, which could increase the cost of generation and emissions from the power system.
This work seeks to explore the potential benefits of active electricity market participation by electrolyzers, considering the viewpoint of both the electrolyzers and the grid. We derive cost curves that can model the bidding behavior of electrolyzers in electricity markets, specifically considering a Renewable Electrolyzer Plant (REP) consisting of a co-located RES and electrolyzer with a common grid connection. Since co-location setup reduces the dependency on the transmission grid and allows for direct consumption of renewable power, it is an economically and operationally attractive configuration for large-scale electrolyzer plants. We then analyze the impact of active market participation of large-scale electrolyzers on electrolyzer profits and power system operations.
I-A Literature review
The market-based scheduling of electrolyzers or REPs is typically modeled through a profit-maximization problem, assuming a day-ahead power price forecast. References [8], [9], [10], and [11] consider different modeling approaches for electrolyzer or REP scheduling, with the focus of accurately representing the physics of the electrolyzer. These papers point to the importance of accurately representing the non-convex hydrogen production curve, which defines the hydrogen produced as a function of the electricity consumed. They perform optimal scheduling under the assumption of an exogenously determined (yet typically time-varying) electricity price, but do not account for the impact of electrolyzer scheduling on electricity network constraints and the electricity price formation. As the electrolyzer capacity expands, the consumption of electrolyzers will increasingly affect grid congestion and cause the electricity price to change. As a result, assuming that the electricity price is an exogenous signal as in [8]-[11] can lead to infeasible or suboptimal schedules, misleading estimates of profits, and negative impacts on the power system operations.
The impact of electrolyzers on power and energy systems has been explored primarily through capacity expansion models. A technology-agnostic study was conducted in [12], employing a generation capacity expansion model to evaluate the role of “demand-sinks”, including, e.g., electrolyzers, in the grid under various cost assumptions. The study finds that demand-sinks might enable additional renewable capacity in the power system while having little effect on the cost of electricity. The demand sinks are assumed to be flexible and are dispatched assuming a fixed utility per consumed energy. Reference [13] explores the impact of various hydrogen investment support policies on the interaction of hydrogen, electricity and emission markets via a partial equilibrium model. Papers [14, 15], and [16] investigate the emissions associated with grid-connected electrolytic hydrogen production, evaluating the effects of proposed regulations such as the temporal matching between electrolyzer consumption and renewable production and the additionality of matched renewable sources. These papers represent the transmission network as either a zonal (i.e., modeling transmission capacities between regions only) or copper-plate (i.e., unconstrained transmission capacities) system. However, [17] argues that not considering the deliverability of renewable energy, i.e., through a zonal or copper-plate representation of the transmission network, drastically changes the analysis of emissions associated with a flexible load, such as a grid-connected electrolyzer.
The literature lacks investigations of the effects of electrolyzers (or REPs) in the power system that accurately represent both the hydrogen production curve of the electrolyzer and models the intra-regional transmission constraints of the power system. Further, there has been limited analysis on the difference in effects when the electrolyzer is operated with a fixed power consumption versus when its implicit flexibility is leveraged through active market participation.
I-B Novelty and contribution
To address the identified gaps in the literature, we make two main contributions. First, we derive the marginal cost curve of REPs (or stand-alone electrolyzers), capturing the non-constant efficiencies of the hydrogen production curve and the time-varying availability of renewable generation. The cost curve is then converted to a bid curve that can be used for active participation in day-ahead and real-time electricity markets, representing either the opportunity cost of selling renewable generation to the grid rather than consuming it for hydrogen production or the revenue derived from importing electricity (if there is insufficient local renewable generation available). Second, we demonstrate how the electrolyzer bid curve can be used to analyze the scheduling of REPs in electricity markets under perfect competition. Specifically, we compare a market-bidding REP, which bids based on the derived cost curve, with a fixed-consumption REP, where the electrolyzer operates at a constant load. Our analysis examines how these strategies affect generation costs, renewable power curtailment, and system emissions, highlighting the influence of the REP’s marginal cost and its placement within the overall system merit order. Finally, we investigate how transmission constraints impact the perceived effects of connecting an electrolyzer to the power system.
(a) Hydrogen production
(b) Hydrogen revenue
(c) REP cost curve
(d) REP marginal cost curve
Figure 1: Derivation of the REPs (marginal) cost curve from the empirical hydrogen production curve. Illustrated for $\bar{P}^{\rm h}=P^{\rm RES}=500$ MW.
II Methodology
This section first presents the derivation of the REPs’ continuous, non-convex, (marginal) opportunity cost curve from an empirical hydrogen production curve. We then approximate this cost curve as a convex, piecewise linear cost curve suitable for electricity market bidding. Finally, we present the mathematical formulation of our assumed electricity market clearing model, a DC Optimal Power Flow (DC OPF) problem.
II-A Derivation of the REP’s Opportunity-Cost Curve
To model the REP as an active electricity market participant, we first derive its cost curve. Typically, renewable generators have zero marginal cost, as there is no variable cost associated with production. However, in a REP that includes both renewable generation and an electrolyzer, the situation is different. While producing hydrogen, the REP derives a revenue from consuming electricity. Selling this electricity to the grid instead would thus be associated with an opportunity cost, leaving the REP with a non-zero cost of exporting electricity. In the following, we derive a cost curve for the REP based on the opportunity cost associated with lost hydrogen production. The cost curve can be used to model the REP’s bids for electricity market-clearing models. From the cost curve, we further derive the REPs marginal cost curve. The steps of the derivation are described in detail below and illustrated in Fig. 1.
II-A1 Hydrogen production curve
The starting point of the derivation is the hydrogen production curve ${\rm h}(p^{\rm h})$ of the electrolyzer [18]. This curve expresses the amount of hydrogen ${\rm h}$ produced per hour at a given power consumption $p^{\rm h}$. Here, we leverage the empirical hydrogen production curve for an alkaline electrolyzer from [8], which is shown in Fig. 1(a). Due to low efficiencies at low and high loads, the hydrogen production curve is non-convex.
II-A2 Hydrogen revenue
Assuming that the produced hydrogen can be sold at a constant price $\lambda^{\rm{h}}$, we derive the hydrogen revenue for a given power consumption ${\rm r}_{\rm h}(p^{\rm h})$ as the product of the hydrogen production and the price of hydrogen,
${\rm r}_{\rm h}(p^{\rm h})={\rm h}(p^{\rm h})\lambda^{\rm{h}}$ $p^{\rm h}\in[0,\bar{P}^{\rm h}],$ (1)
where $\bar{P}^{\rm h}$ is the electrolyzer capacity. The resulting hydrogen revenue curve is plotted in Fig. 1(b).
For an electrolyzer within a REP where the renewable generation capacity is larger than the electrolyzer capacity, i.e., $P^{\rm RES}>\bar{P}^{\rm h}$, the available renewable power may at times exceed the electrolyzer capacity. In these cases, the excess power cannot be used to further increase hydrogen production. The maximum hydrogen revenue is therefore equal to the revenue of maximum hydrogen production, ${\rm r}_{\rm h}(\bar{P}^{\rm h})$. To reflect this case, we formulate a piece-wise curve, defining the hydrogen revenue for any available (positive) power:
${\rm r}_{\rm h}(p^{\rm h})=\begin{cases}{\rm h}(p^{\rm h})\lambda^{\rm{h}},&p^{\rm h}\in[0,\bar{P}^{\rm h}],\\ {\rm h}(\bar{P}^{\rm h})\lambda^{\rm{h}},&p^{\rm h}>\bar{P}^{\rm h}.\end{cases}$ (2)
The first piece follows what is illustrated in Fig. 1(a), and the second piece is a constant equal to the revenue of maximum hydrogen production.
II-A3 Opportunity cost
For the REP to sell power rather than produce hydrogen, the revenue from selling power must exceed the revenue that the REP would get from withholding power for hydrogen production. The REP’s cost of selling a power $p^{\rm DA}$ is thus an opportunity cost, equal to the lost revenue of not producing hydrogen from the sold power.
To obtain the opportunity cost, we first derive the revenue the REP can achieve by withholding all the available renewable power $P^{\rm RES}$. We insert $P^{\rm RES}$ into (2) and obtain ${\rm r}_{\rm h}(P^{\rm RES})$, which for a given available renewable power is a constant. Following ${\rm r}_{\rm h}$’s piece-wise definition, ${\rm r}_{\rm h}(P^{\rm RES})$ is capped at the revenue of maximum hydrogen production if $P^{\rm RES}$ exceeds the electrolyzer capacity $\bar{P}^{\rm h}$.
Next, we look at the revenue achieved from hydrogen production as a function of power sold to the DA market. When selling a power $p^{\rm DA}$, the power available for hydrogen production is given by $p^{\rm h}=P^{\rm RES}-p^{\rm DA}$. Inserting into (2), we achieve ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$. Again, this revenue is capped at the revenue associated with maximum hydrogen production and takes this constant value when $P^{\rm RES}-p^{\rm DA}>\bar{P}^{\rm h}$.
The opportunity cost ${\rm c}_{\rm el}(p^{\rm DA})$ of selling a power $p^{\rm DA}$ is equal to the difference between the hydrogen revenue of withholding all available renewable power and the hydrogen revenue when selling a power $p^{\rm DA}$,
${\rm c}_{\rm el}(p^{\rm DA})$ $={\rm r}_{\rm h}(P^{\rm RES})-{\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA}),$ (3)
where $p^{\rm DA}$’s domain depends on the available renewable power relative to the electrolyzer capacity,
$p^{\rm DA}\in\begin{cases}[P^{\rm RES}-\bar{P}^{\rm h},P^{\rm RES}],&P^{\rm RES}\leq\bar{P}^{\rm h},\\ [0,P^{\rm RES}],&P^{\rm RES}>\bar{P}^{\rm h}.\end{cases}$ (4)
Following from eq. (4), the REP can always export its full available renewable capacity, but can also bid to import power, i.e., have a negative $p^{\rm DA}$, if its available renewable production $P^{\rm RES}$ is lower than the electrolyzer capacity $\bar{P}^{\rm h}$. In reality, a negative $p^{\rm DA}$ represents a willingness-to-pay to import power, and not a cost. For simplicity, we still refer to the full curve as the (opportunity) cost curve.
The cost curve, plotted for the case where $P^{\rm RES}=\bar{P}^{\rm h}$ in Fig. 1(c), is a shifted reflection of the hydrogen revenue curve over both axes. First, as the cost curve, a function of $p^{\rm DA}$, is a constant ${\rm r}_{\rm h}(P^{\rm RES})$ subtracted by the revenue function ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$, its resulting curve is a vertical reflection of the hydrogen revenue curve, shifted along the y-axis with the value of ${\rm r}_{\rm h}(P^{\rm RES})$. Further, as the variable $p^{\rm DA}$ is negatively represented in the hydrogen revenue function, ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$, it is also a horizontal reflection of the hydrogen revenue curve, shifted along the x-axis with the constant $P^{\rm RES}$. This implies that as the hydrogen revenue curve is (approximately) concave, the REP’s cost curve is (approximately) convex.
In the cases where $P^{\rm RES}$ exceeds the electrolyzer capacity, ${\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$ is equal to the maximum hydrogen revenue ${\rm r}_{\rm h}(P^{\rm RES})$ for $p^{\rm DA}\in[0,P^{\rm RES}-\bar{P}^{\rm h}]$. For this range of sold power, the opportunity cost function thus evaluates to zero, indicating that there is no opportunity cost associated with selling renewable power that exceeds what the electrolyzer can consume. The cost curve then forms a piece-wise function, as follows from eq. (2).
The presented method allows for computing the cost curve of a stand-alone electrolyzer, by setting $P^{\rm RES}$ to zero.
Figure 2: Illustrations of the REP’s approximated cost curves (first row) and approximated marginal cost curves (second row). We show three cases of renewable power production relative to the electrolyzer capacity, set to 500 MW in this example, with four arbitrarily selected pieces. When there is excess renewable power, meaning production exceeds the electrolyzer capacity, this excess power is offered at zero marginal cost, as it cannot be used for hydrogen production and has no associated opportunity cost. The negative values of the cost curve to the right indicate that the REP bids to purchase the corresponding quantity of power, at a price up to the marginal cost for that quantity.
II-A4 Marginal opportunity cost
We find the REPs marginal cost of selling power ${\rm mc}_{\rm el}(p^{\rm DA})$ as the change in opportunity cost $d{\rm c}_{\rm el}$ over the change in sold power $dp^{\rm DA}$, i.e., the derivative of the opportunity cost,
${\rm mc}_{\rm el}(p^{\rm DA})={\rm c}_{\rm el}^{\prime}(p^{\rm DA})=d{\rm c(p^{\rm DA})}/dp^{\rm DA}.$ (5)
The marginal cost curve of the REP in the case where $P^{\rm RES}=\bar{P}^{\rm h}$ is illustrated in Fig. 1(d).
II-B Piecewise-linear approximation of the cost curve
As seen in Figs. 1(c),1(d), the REP (marginal) cost curve is non-convex. However, most electricity markets require participants to provide piecewise linear, convex cost curves, which corresponds to piecewise constant, non-decreasing marginal costs. We therefore derive an approximated cost curve.
Instead of directly approximating the opportunity cost curve in Fig. 1(c), we begin by approximating the hydrogen production curve in Fig. 1(a). As will become evident, this allows us to express our approximated opportunity cost curve directly as a function of the hydrogen price $\lambda^{\rm{h}}$, which allows for easier modification in case of hydrogen price changes. We approximate the hydrogen production curve as a concave piecewise linear function with decreasing slope:
$h(p^{\rm{h}})=\sum_{i\in\mathcal{I}}\left(A_{i}p^{\rm{h}}+B_{i}\right)\mathbf{1}_{\left(p^{\rm{h}}\in[\underline{P}_{i},\overline{P}_{i}]\right)},$ (6)
where $A_{i}$ and $B_{i}$ represent the slope and constant corresponding to the $i^{th}$ linear piece, and $[\underline{P}_{i},\overline{P}_{i}]$ is the range of electrolyzer power consumption corresponding to piece $i$. The set $\mathcal{I}$ represents the set of pieces, and $\mathbf{1}_{(.)}$ is the indicator function which takes on a value of 1 when $p^{\rm{h}}\in[\underline{P}_{i},\overline{P}_{i}]$ (and zero otherwise). The approximation is chosen such that the slope is decreasing for all pieces, i.e., $A_{i}>A_{i+1}$.
From the approximated hydrogen production curve we derive the approximated opportunity cost for selling $p^{\rm DA}$:
${\rm c}_{\rm el}(p^{\rm{DA}})={\rm r}_{\rm h}(P^{\rm RES})-{\rm r}_{\rm h}(P^{\rm RES}-p^{\rm DA})$
$=\lambda^{\rm h}\sum_{i\in\mathcal{I}}\left(A_{i}P^{\rm RES}+B_{i}\right)\mathbf{1}_{\left(P^{\rm{RES}}\in[\overline{P}_{i},\underline{P}_{i}]\right)}$ (7)
$-\lambda^{\rm h}\sum_{i\in\mathcal{I}}\left(A_{i}(P^{\rm RES}-p^{\rm DA})+B_{i}\right)\mathbf{1}_{\left((P^{\rm{RES}}-p^{\rm{DA}})\in[\overline{P}_{i},\underline{P}_{i}]\right)}.$
Since we approximated the hydrogen production curve as a concave, piecewise linear function, the resulting cost curve forms a convex, piecewise linear function, with increasing slopes. It can therefore be directly implemented into, for example, a DC OPF, without the need for binary variables.
The marginal opportunity cost is the derivative of the opportunity cost, which for the piece-wise linear curve is the slope of each piece with respect to $p^{\rm DA}$. The approximated marginal cost therefore becomes an increasing, step-wise function:
${\rm mc}_{\rm el}(p^{\rm{DA}})=\sum_{i\in\mathcal{I}}\lambda^{\rm h}A_{i}\mathbf{1}_{\left((P^{\rm{RES}}-p^{\rm{DA}})\in[\overline{P}_{i},\underline{P}_{i}]\right)}.$ (8)
Depending on the electrolyzer capacity relative to the available renewable power, three different cases of the (marginal) cost curve arise, as illustrated in Fig. 2:
  • If the renewable power production is equal to the electrolyzer capacity, $P^{\rm RES}=\bar{P}^{\rm h}$, any sold power is associated with a (non-zero) positive opportunity cost, forming a cost curve that intercepts with the origin. This case is illustrated by the first column of plots in Fig. 2, and corresponds to the approximation of the curves in Fig. 1.
  • If there is excess renewable power production, $P^{\rm RES}>\bar{P}^{\rm h}$, the cost curves are shifted right to higher values of $p^{\rm DA}$. For the excess renewable generation, we define an initial piece with $\underline{P}_{1}=0,\overline{P}_{1}=P^{\rm{RES}}-\overline{P}^{\rm h}$, for which the cost is zero, i.e., $A_{1}=b_{1}=0$. This case is illustrated in the second column of graphs in Fig. 2.
  • If there is insufficient renewable power, $P^{\rm RES}<\bar{P}^{\rm h}$, the marginal cost curve is shifted left towards lower values of $p^{\rm DA}$, and down towards lower costs. The negative bid curve at negative values of $p^{\rm DA}$ is interpreted as the marginal bid for “negative exports”, which corresponds to imports, i.e., consumption, of electricity. This case is illustrated by the plots in the last column of Fig. 2.
II-C Integrating Electrolyzers in the Electricity Market Clearing
The market-bidding REP submits its approximated cost curve, derived in the previous section, to the system operator and receives a generation schedule $p^{\rm DA}$ after the electricity market has cleared. The REP then uses the remaining power for hydrogen production (up to the full capacity of the electrolyzer), i.e., $p^{\rm h}=\min\left(P^{\rm RES}-p^{\rm DA},\bar{P}^{\rm h}\right)$.
To model the electricity market clearing, we solve a standard DC OPF. We assume a system with a set of $\mathcal{N}$ nodes, indexed by $n$. For simplicity of notation, but without loss of generality, we assume that we have one generator $p_{n}^{\rm G}$ and one inflexible demand $P_{n}^{\rm D}$ per node. The cost of generator $n$ is modeled using piecewise linear costs curves with a set of $\mathcal{I}_{n}$ pieces and the auxiliary variable $c_{n}^{\rm G}$ representing the total cost. With this notation, we formulate the DC OPF as follows:
$\allowdisplaybreaks\underset{c_{n}^{\rm G},p_{i}^{\rm G},\theta_{i}}{\operatorname{min}}$ $\sum_{n\in\mathcal{N}}c_{n}^{\rm G}$ (9a)
s.t. $c_{n}^{\rm G}\geq\alpha_{i}p_{n}^{\rm G}+\beta_{i}$ $\!\!\forall i\in\mathcal{I}_{n},n\in\mathcal{N}$ (9b)
$p_{n}^{\rm G}\!-\!P_{n}^{\rm D}=\!\!\!\!\!\sum_{m\in\mathcal{K}_{n}}\!\frac{1}{X_{nm}}\left(\theta_{n}-\theta_{m}\right)$ $\!\!\forall n\in\mathcal{N}~{}~{}:\lambda^{\rm DA}_{n}$ (9c)
$\!\!\!\!\!-\!\bar{P}_{nm}\leq\!\frac{1}{X_{nm}}\left(\theta_{n}-\theta_{m}\right)\leq\bar{P}_{nm}$ $\!\!\forall m\in\mathcal{K}_{n},n\in\mathcal{N}$ (9d)
$\underline{P}_{n}^{\rm G}\leq p_{n}^{\rm G}\leq\bar{P}_{n}^{\rm G}$ $\!\!\forall n\in\mathcal{N}$ (9e)
$\theta_{n=1}=0,$ (9f)
where the variables $c_{n}^{\rm G}$, $p_{n}^{\rm G}$ and $\theta_{n}$ denote the total generation cost, power generation schedule, and voltage angle at node $n$, respectively. The objective function (9a) minimizes the total cost of generation. Constraint (9b) implements the piecewise linear, convex cost curves with parameters $\alpha_{i}$ and $\beta_{i}$, representing the slope and intercepts for each piece. Note that the inequality will always be binding (i.e., satisfied with equality) as the cost variables $c_{n}^{\rm G}$ are minimized. The reactance of each branch is given by ${X_{nm}}$, where $m\in\mathcal{K}_{n}$ notes all nodes $m$ connected to $n$. Constraint (9c) enforces the nodal power balance, while (9d) and (9e) ensure minimum and maximum limits on branch power flow and generation capacity. Finally, (9f) defines the reference node. The Locational Marginal Prices (LMPs) in the day-ahead stage are given by the dual variable associated with the power balance constraints (9c), denoted as $\lambda^{\rm DA}_{n}$. REPs are modeled as generators in this formulation, with the day-ahead schedule $p^{\rm DA}$ represented by a generator variable $p_{n}^{\rm G}$. Here, we use the slope and intercepts as derived in (II-B) to represent the REP. However, assuming no start-up costs, the approximated marginal cost curve is equivalent to the full cost curve and can also be implemented in a DC-OPF. Further, the marginal cost curve is important for understanding the REP’s position in the merit order curve.
Note that cases where the REP imports power are accommodated by adapting the generation limits to allow for negative generation. We consider no inter-temporal constraints such as ramping rates or start-up time.
III Case study
This section analyzes the impact of connecting an electrolyzer to the grid, when co-located with a renewable power plant. We compare active market participation through the proposed bidding approach to the traditional approach of fixed electrolyzer consumption.
III-A Case Study Description
Figure 3: Generation capacities and demand range per region of the RTS-GMLC. While region 3 has the most capacity, it also has a slightly higher demand load, approximately 10% higher than the other two regions across the year. Further, the increase in generation capacity in region 3 is dominated by renewable power sources. For conventional sources, the generation capacity is similar in each region.
We define our case study based on the Reliability Test System—Grid Modernization Lab Consortium (RTS-GMLC) [19], an updated version of the IEEE RTS-96 with high shares of renewable power capacity. The RTS-GMLC system is divided into three regions, and comes with one year of data for hourly load and renewable production, as well as piece-wise cost data for all conventional generators. The load in the system is inflexible. We refer to the RTS-GMLC system without any modifications, i.e., without any electrolyzer capacity, as the base system. The production capacities and demand range of each region in the system are plotted in Fig. 3. Note that the majority of the renewable capacity in the system is located in region 3.
Figure 4: Merit order curve of the total installed generation capacities for a low (left) and high (right) hydrogen price. The REP illustrated consists of a 847 MW wind farm and a 500 MW electrolyzer capacity. The adjusted net demand, i.e., the demand subtracted by the renewable production of every time-step, adjusted by adding the total renewable capacity in the system, is indicated by the shaded area. When minimizing cost of generation, only the generation to the left of the adjusted net demand will be dispatched, unless there is congestion in the network.
We select a wind farm, generator no. 156 located at node 303 in region 3, to convert to a REP. This wind farm has a total installed capacity of 847 MW and an average production of 237 MW throughout the year. We consider three different electrolyzer capacities, namely 100 MW, 500 MW and 1000 MW, but keep the REP’s renewable capacity equal to the pre-existing capacity in all cases. Hence, the only change to the base system is the added electrolyzer capacity at node 303.
We consider two different hydrogen prices, \$1.5 and \$6 per kg. These prices are representative of the current market rate for conventionally produced hydrogen and the current market rate for green hydrogen with a subsidy [2], respectively. For every hour, we calculate the marginal cost curve of the REP based on the production of the local wind farm, the capacity of the electrolyzer, and the assumed hydrogen price.
III-B Analysis of the Merit Order Curve
To provide an insight into how the REP’s marginal cost of electricity generation compares with other generators in the system, we first analyze where the REP falls in the merit order. Fig. 4 shows the merit order curve of the generation stack of the entire RTS-GMLS system for the two different hydrogen prices, both with an electrolyzer capacity at the REP of 500 MW. The range of adjusted net-demand is marked by the purple shaded area. The adjusted net-demand of an hour is the total demand minus the total renewable production, adjusted by adding the total renewable capacity of the system. The range indicates all possible values of adjusted net demand throughout the year. In the absence of congestion, generators that fall in this range would be the marginal generator at certain hours. Generation capacity that falls to the right of the adjusted net demand range will only be dispatched in the event of network congestion.
For both the lower and higher hydrogen prices, the REP (marked in red) has a piece of its marginal cost curve equal to zero, as its renewable capacity of 847 MW exceeds its electrolyzer capacity of 500 MW. However, the presence and size of the zero marginal cost piece will depend on the renewable power availability of every hour. For the lower hydrogen price (left plot), the remaining pieces of the REP cost curve have marginal cost similar to that of natural gas generation and fall partly within the adjusted net demand range. For the higher hydrogen price (right plot), the remaining pieces of the REP cost curve have higher marginal costs, similar to that of oil-fueled generation. These pieces fall outside the adjusted net demand range, indicating that this power will only be exported when network congestion causes prices to increase.
III-C Simulation and Benchmarking Framework
To analyze the impact of a market bidding REP on the system and benchmark its effects against a more traditional style of operation, we simulate the electricity market clearing by solving the DC OPF problem (9) for each hour of the year. To solve the DC OPF, we use the implementation provided in the Julia package PowerModels.jl [20]. Our source codes are publicly available in [21].
III-C1 REP Representations
We compare our proposed marginal cost bidding approach with a more traditional fixed consumption case. Below, we describe each case in detail:
Market bidding REP: The REP participates in the market, submitting its opportunity cost curve. A DC OPF is solved for the RTS-GMLC with the cost curve of wind farm 156 altered to represent the presence of the electrolyzer as described in Section II-C. This case represents how the REP would behave when participating in the electricity market and fully exposed to the LMPs of the system (i.e., fully flexible).
Fixed consumption REP: We compare the market bidding REP to an REP where the electrolyzer is consuming at a fixed, constant power. To create a fair comparison, we define the fixed consumption power as the average consumption of the market bidding REP, i.e., the total electricity consumption of the market bidding REP over the year, divided by the number of hours in the year. The electrolyzer is modeled as an additional load at the REP’s node, not modifying the bid curve of wind farm 156. This new load is set to the calculated fixed consumption and is inflexible. However, for several hours, this level of consumption causes the DC OPF to become infeasible. In those hours, we first model the fixed consumption as a bid with a high marginal cost (of \$10,000 per MWh), thus approximating the highest electrolyzer consumption possible while maintaining a feasible program. We then rerun the DC OPF with the new load set equal to this (lower) consumption.
III-D System-focused Results
Figure 5: Total cost of generation and cost of generation per load for a fixed consumption and market bidding electrolyzer under the hydrogen prices of \$currency-dollar\\$\$1.5 and \$currency-dollar\\$\$6 per kg.
We first investigate the impact of the REP’s electrolyzer from a system perspective, considering the cost of generation, renewable power curtailment, and emissions. We combine the three electrolyzer capacities (100, 500 and 1000 MW), two hydrogen prices (\$1.5 and \$6 per kg) and two REP representations previously described, resulting in 12 cases.
III-D1 Cost of generation
We examine the average cost of generation throughout the year for each case. Due to the REP’s ability to act as a consumer, the objective value of the DC OPF consists not only of the cost of generation, but also of the utility of consumption from the REP. To evaluate the changes in cost of generation due to the presence of the electrolyzer, we compute the cost of generation of each hour as:
Cost of $\text{Cost of generation}=\sum_{g\in\mathcal{G\backslash\text{REP}}}c_{g}^{\rm G},$ (10)
where $\mathcal{G\backslash\text{REP}}$ is the set of all generators apart from the REP. The cost of generation is therefore not equal to the objective value (9a). We sum across all hours to achieve the total cost of generation, and divide by the total generation in the system to obtain the cost of generation per load. The results are shown in Fig. 5, which also reports changes in cost relative to the base system.
We observe that the total cost of generation (two left plots) increases as the electrolyzer capacity increases. This is expected as the electrolyzer imposes an additional load on the system. For the smaller electrolyzer capacity of 100 MW, the cost increases by less than 2.5% for any operational mode or hydrogen price. However, for larger electrolyzer capacities, the total cost of generation increases by 10-20%. In general, the market bidding REP yields a smaller increase in the total cost of generation compared to the corresponding fixed consumption case. This is expected as the market bidding REP will reduce its consumption if the LMP is sufficiently high. However, for an electrolyzer capacity of 1000 MW operating at the higher hydrogen price, the market bidding REP causes a higher cost of generation than the fixed consumption counterpart. This is because the fixed consumption REP is consuming significantly less power overall due to congestion in the network, as will be elaborated later in Section III-E1.
Next, we consider the cost of generation per load, shown in the two plots to the right in Fig. 5. We first observe that adding a small electrolyzer with 100 MW capacity decreases the cost of generation per load for both hydrogen prices and REP representations. This is because part of added REP consumption is met by otherwise curtailed renewable power. For electrolyzer capacities of 500 and 1000 MW, the average cost generally increases. However, the magnitude of the increase depends on the case. Cases with fixed consumption and/or a higher hydrogen price lead to significant increases in average cost, reaching up to 7. 7% (although a market bidding REP causes smaller increases compared to a fixed consumption REP). In contrast, adding a market bidding REP with a lower hydrogen price yields a negligible increase in generation cost per load.
From these results, we conclude that to keep the cost of generation low, the REP must be sensitive to the LMP, i.e., operated as a market bidder. Further, higher hydrogen prices reduce the benefits of market bidding REP operation. Following the merit order curve in Fig. 4, the market bidding REP with a high marginal cost is only dispatched if there is network congestion, and will therefore impact the total cost of generation in a way similar to that of a fixed consumption REP. We observe this effect in the cases with a hydrogen price of \$6 per kg and large electrolyzer capacities, and note that subsidies for hydrogen production, through inflating the hydrogen price, might result in less flexible operation of the REP and thus increase the cost of generation of its connected system.
Figure 6: Total curtailment of renewable power generation.
III-D2 Renewable power curtailment
We next investigate how the presence of an electrolyzer in the system affects the curtailment of renewable power, as it is expected that electrolyzers can use surplus renewable power and thus reduce the need for curtailment. Fig. 6 shows the total renewable power curtailment over the year in the entire system, for the different electrolyzer capacities, REP representations, and hydrogen prices explored previously. Intuitively, as the load in the system increases, the amount of curtailed renewable power should decrease. This trend is observed in the market bidding cases with a low hydrogen price. However, in the other cases, increasing the electrolyzer capacity from 500 to 1000 MW actually increases the curtailment. This increase in curtailment is a result of congestion in the network. When the REP bids to import power at a sufficiently high price (or in the fixed consumption cases), congestion forces the DC OPF to dispatch expensive, non-renewable generation and curtails renewable generation to allow the REP to import more power.
We draw a conclusion similar to that in Section III-D1, that high hydrogen prices reduce the flexibility of electrolyzers, potentially leading to negative impacts on RES curtailment.
III-D3 System emissions
Finally, we investigate the emissions associated with the generation of power in the system. We compute the total emissions and emissions per load, i.e., total emissions divided by total system load across all hours, as well as the increase from the base system for the 12 described cases. The results, reported in Table I, are given in CO2 equivalents. We assume constant emission factors of 0.9606, 0.6042, and 0.7434 metric tons of CO2 per MWh for coal, natural gas, and oil-fueled power plants, respectively, and zero emissions for renewable and nuclear generation.
Our results show that for a given electrolyzer capacity, both the total emissions and emissions per load remain similar irrespective of both how the REP is operated and the hydrogen price. This result is understood by examining the merit order curve of the system in Fig. 4. The cheapest generation sources, i.e., renewable and nuclear capacities, are indeed also the least emission intense. However, the (generally) cheapest conventional generation capacities, the coal-fueled plants, have higher emission factors than the more expensive natural gas and oil-fueled plants. Therefore, following price-signals cannot guarantee lower system emissions.
The emissions generally increase with the installed electrolyzer capacity. For the 100 MW electrolyzer, this increase is negligible, as about 20% of the additional load the electrolyzer imposes on the system is met with otherwise curtailed renewable power. For larger electrolyzer capacities, conventional generation meets a larger share of the additional system load, resulting in increased system emissions.
We conclude that market bidding is only an efficient measure to avoid increases in system emissions under strong correlation between the marginal cost and emissions of generation.
TABLE I: System emissions results over the year.
Total emissions Emissions per load
Case [M t] $\Delta$ [kg/MWh] $\Delta$
Base system 13.88 - 368.50 -
100 MW \$currency-dollar\\$\$1.5/kg bidder 14.38 3.61% 373.83 1.45%
fixed 14.38 3.62% 373.77 1.43%
\$currency-dollar\\$\$6/kg bidder 14.43 3.97% 374.39 1.60%
fixed 14.43 3.97% 374.39 1.60%
500 MW \$currency-dollar\\$\$1.5/kg bidder 16.39 18.13% 396.37 7.56%
fixed 16.39 18.10% 396.15 7.50%
\$currency-dollar\\$\$6/kg bidder 16.88 21.62% 401.42 8.93%
fixed 16.88 21.64% 401.50 8.96%
1000 MW \$currency-dollar\\$\$1.5/kg bidder 17.70 27.54% 410.67 11.44%
fixed 17.69 27.46% 409.53 11.13%
\$currency-dollar\\$\$6/kg bidder 18.08 30.29% 414.96 12.61%
fixed 18.32 32.00% 415.68 12.80%
III-E REP-focused results
TABLE II: REP results over the year.
Electrolyzer REP Average
average load profit LMPa
Case [MW] [M \$] [\$/MWh]
100 MW \$$\$1.5$/kg bidder 91.44 28.88 17.81
fixed 91.44 28.14 17.75
\$$\$6$/kg bidder 100.00 98.12 17.83
fixed 100.00 98.12 17.83
500 MW \$$\$1.5$/kg bidder 421.15 59.73 21.55
fixed 421.15 47.36 24.30
\$$\$6$/kg bidder 499.12 362.15 31.77
fixed 499.01 352.36 33.92
1000 MW \$$\$1.5$/kg bidder 629.77 67.76 24.45
fixed 617.38 -730.61 179.60
\$$\$6$/kg bidder 729.34 320.83 81.54
fixed 672.3 -563.48 241.06
a Average LMP at the REP’s node.
We now analyze our results from the perspective of the REP. For each of the 12 cases described previously, we calculate the average load of the electrolyzer, the REP’s total profit and the average LMP at the REP’s node. The total profit is calculated considering the electricity cost and revenue (assuming that the REP is remunerated for exports or pays for imports based on the hourly nodal LMP) as well as the revenue from hydrogen production (using the hydrogen price for the specific case). The results for the 12 REP cases are provided in Table II.
III-E1 Electrolyzer utilization
We first discuss the average load of the electrolyzer in the various cases, shown in the first column of Table II. For the cases of the higher hydrogen price and installed electrolyzer capacity of 100 MW and 500 MW, the average load is 100 MW and $>$499 MW, indicating that the overall utilization is (near) 100%. For the same electrolyzer capacities, but at the lower hydrogen price, the average load is reduced to an overall utilization of 91% and 84%. These results align well with our analysis of the merit order curve in Fig. 4. For a higher hydrogen price, the REP lies high in the merit order curve, and it would seldom be beneficial to reduce consumption to sell electricity. For a lower hydrogen price, the marginal opportunity cost is lower, and selling electricity instead of producing hydrogen is cost-optimal in more hours. Similarly, the lower marginal revenue associated with a lower hydrogen price would cause the REP to import less electricity.
For an electrolyzer with 1000 MW capacity, we observe that the average load is significantly below capacity for the market bidding REPs, even for the case with a high hydrogen price. This indicates that a combination of congestion and high prices at the REP node reduces the viability of importing sufficient power to run the electrolyzer at full capacity.
Further, while the fixed consumption REP should consume the average electrolyzer load of the corresponding market bidding case, we observe that the actual average consumption is significantly lower than for the market bidding REP. As previously explained, the fixed consumption level leads to an infeasible program in certain hours, and for those hours we reduce the consumption level of the fixed electrolyzer. The lower average load therefore indicates that a fixed consumption at this level is frequently infeasible.
III-E2 Profit
We next discuss the REP’s operational profit under the different cases, shown in the second to last column of Table II. Recall that the REP profit is calculated assuming that the REP imports and exports power at a price equal to the LMP for all cases. Under this assumption, we hypothesize that the market bidding REP will achieve a profit equal to or higher than that of the fixed consumption REP, since the market bidding REP is actively adapted to the LMP. From our results, we observe that this hypothesis holds true. For the smaller electrolyzer capacity and high hydrogen price, the profit is similar for market bidding and fixed consumption operation, as it is cost-effective to run the electrolyzer at full capacity in both cases. However, if the hydrogen price is lower and/or the electrolyzer capacity is sufficiently large (e.g., 500 or 1000 MW) to cause congestion or infeasibility, we observe that there can be significant differences in profit between the market bidding and fixed consumption cases. In the case of a 500 MW electrolyzer capacity and a lower hydrogen price, the market bidding case yields a 26% profit increase compared to the fixed consumption case. In the case of a 1000 MW electrolyzer capacity, the profits achieved through active market bidding far exceed those of a fixed consumption, which both exhibit a loss (i.e., a negative profit).
Finally, we observe that a larger electrolyzer is not always better in terms of profit. When we compare profits for market bidding REPs with 500 and 1000 MW electrolyzer capacity, we observe that a larger capacity increases profits in the case with a \$1.5 per kg hydrogen price but decreases profits in the case with a \$6 per kg hydrogen price. The average LMP at the REP’s node, given by the last column of Table II, shows that there is a significant increase in LMP when the electrolyzer capacity increases from 500 to 1000 MW in certain cases. We examine the case of a market bidding REP under the higher hydrogen price. In this case, the LMP increases from \$31.77 to \$81.54 per MWh. Recalling the merit order curve of Fig. 4, the marginal cost of the REP under a higher hydrogen price lies in the range of \$80-120 per MWh. The REP therefore makes more profit from producing hydrogen than from selling electricity at LMPs up to this range. However, with increasing LMP, the profit associated with hydrogen production decreases. The increase in electrolyzer capacity when operating under the higher hydrogen price increases the LMP to a point where the REPs profit decreases.
We conclude that to evaluate the profit of a market bidding REP, it could be essential to model the on the price formation process through, e.g., a DC OPF, as the presence of the electrolyzer in the system can highly impact the LMPs and the REP’s profit. This becomes more important with larger electrolyzer capacities and higher hydrogen prices.
TABLE III: System-focused results under nodal, zonal, and copper-plate modeling. The results are reported for the base system and two market bidding REPs, both with 1000 MW electrolyzer installed.
System Cost of System Renewable
load generation emissions curtailment
Case [TWh] [\$currency-dollar\\$\$/MWh] [t/MWh] [GWh]
Base Copper 37.66 18.54 0.36 348.75
Zonal 37.66 18.55 0.36 369.97
Nodal 37.66 18.67 0.37 608.09
\$currency-dollar\\$\$1.5/kg Copper 45.07 18.81 0.43 40.47
Zonal 44.94 18.80 0.42 40.77
Nodal 43.19 18.72 0.41 67.29
\$currency-dollar\\$\$6/kg Copper 46.44 19.00 0.43 40.47
Zonal 46.44 19.04 0.43 40.77
Nodal 44.06 19.25 0.42 302.07
III-F The Effect of Transmission Network Modeling
Finally, we investigate the effect of transmission network modeling on the system- and REP-focused results. We compare three network models with varying levels of detail regarding transmission network constraints: a nodal model (which includes all transmission lines), a zonal model (which relaxes all intra-regional transmission constraints), and a copper-plate model (which relaxes all transmission constraints). To create the zonal model, we define each of the three regions in the RTS-GMLC system, as described by Fig. 3, as individual zones. The total inter-regional transmission capacities between regions 1$\leftrightarrow$2, 1$\leftrightarrow$3, and 2$\leftrightarrow$3, are 1175, 500, and 500 MW, respectively. We solve the DC OPF problem for the system with (i) no REP (base case) and (ii) a market bidding, 1000 MW electrolyzer REP with low hydrogen price, and (iii) a market bidding, 1000 MW electrolyzer REP with high hydrogen price. These three cases are solved for each transmission network model, resulting in a total of nine cases. For each case, we compute the system load, cost of generation per load, system emissions per load, and total renewable power curtailment. The system-focused results are reported in Table III. We further report REP-focused results including the average electrolyzer load, REP’s profit, and the average LMP at the REP’s node in Table IV. All results are computed in the same manner as described in Section III-D.
III-F1 System-focused results
We first examine the system load (first column of Table III). For both hydrogen prices, the increase in system load from the base system is lower under a nodal model than under a zonal or copper-plate model. In the case of the higher hydrogen price, the consumption of the REP increases the system load by 23% under a zonal model and by 17% under a nodal model. This is expected, as our analysis in Section III-E showed that the 1000 MW electrolyzer utilization is limited by system congestion and increased LMPs under the nodal model. However, there is no substantial further increase in system load when going from a zonal model to a copper-plate model at any hydrogen price. This suggests that congestion is primarily intra-regional and can only be fully captured by a nodal model.
Next, we discuss the cost of generation per load (second column). We observe that this cost remains relatively unchanged across the different transmission grid models, even though the nodal model serves less load. This indicates that the average cost of generation is higher for the nodal model. This is expected, as congestion in the network, which cannot be captured by a copper-plate model, may lead to the dispatch of more expensive generation in the nodal model.
The system emissions per load (third column) increase when the transmission network modeling level is relaxed. As seen before, large increases in system load is mainly met by conventional generation, which decreases the share of load met by renewable generation in the system. The system emissions per load are therefore higher for the relaxed transmission network models as they allow for higher REP consumption.
The last column of Table III shows the renewable power generation curtailment. These results indicate that the copper-plate and zonal models have much lower curtailment values compared to the nodal model. This is in part because network congestion limits REP consumption and in part because relaxed grid models overestimate the ability to dispatch renewable generation. The copper-plate and zonal network models significantly underestimate the amount of renewable curtailment for the base system and fail to account for the fact that the REP, with a high hydrogen price, increases congestion and subsequent renewable power curtailment.
We conclude that the choice of transmission network model significantly affects the perceived results of adding an electrolyzer to a power system, as the added load could induce congestion in the system.
TABLE IV: REP-focused results under nodal, zonal, and copper-plate modeling. Results are reported for a 1000 MW electrolyzer, market bidding REP, under the lower and higher hydrogen price.
Electrolyzer REP Average
average load profit LMP
Case [MW] [M \$currency-dollar\\$\$] [\$/MWh]
\$currency-dollar\\$\$1.5/kg Copper 843.53 80.92 22.05
Zonal 829.25 79.80 22.26
Nodal 629.77 67.76 24.45
\$currency-dollar\\$\$6/kg Copper 1000.00 769.76 22.39
Zonal 1000.00 766.06 22.85
Nodal 729.34 320.83 81.55
III-F2 REP-focused results
The results for electrolyzer utilization (first column of Table IV) further illustrate to what extent the relaxed grid models overestimate the power consumption of the REP. For example, in the case with a higher hydrogen price, the electrolyzer utilization is at 100% under the relaxed network models, compared to 73% under the nodal model. This increase in utilization contributes to an overestimation of the REP’s profits, which are reported in the second column. Further, the increase in perceived profit is in part due to the LMPs at the REP’s node (last column of Table III). The copper-plate and zonal models yield lower average LMPs, as reduced congestion allows for the dispatch of cheaper generators. In the case of the higher hydrogen price, this leads to vastly differing results on REP profitability. Specifically, the zonal and copper-plate models result in more than twice the REP’s profit compared to the nodal model. The results presented for a nodal model in Section III-E show that a 1000 MW electrolyzer would actually yield a lower profit than a 500 MW electrolyzer in the case of a high hydrogen price, due to the increase in LMPs. This could dis-incentivize larger electrolyzer capacity investments under a nodal pricing scheme. However, the results of profits in Table IV show that this incentive is lost under the zonal and copper-plate models.
We conclude that relaxed transmission models might result in misleading estimates of the REP’s profit as they cannot capture the full effects of congestion on LMPs. Further, these effects lead to investment signals that would be lost under zonal pricing schemes.
IV Conclusions
This paper proposes an opportunity-cost curve for REPs, with the aim of facilitating the integration of such plants in the electricity market. The opportunity cost is derived from the revenue of hydrogen production and forms a curve that can be used to model REP bids in an electricity market. We compare a market bidding REP, bidding its derived opportunity-cost curve, to a REP with a fixed electrolyzer consumption. The results are obtained by solving a DC OPF of the RTS-GMLC test system, modifying a pre-existing wind farm cost curve to represent the REP.
The results show that active market participation can efficiently mitigate increases in the cost of generation per load that occur when the electrolyzer is operated in a fixed manner. Similarly, a market bidding REP efficiently reduces the curtailed renewable power in the system. However, high hydrogen prices will move the REP further up on the merit-order curve, causing it to behave more like a fixed load in the system. Further, market bidding in itself is not necessarily an efficient measure to prevent increases in system emissions due to the lack of correlation between the emission intensities and the marginal cost of generators.
The size of the electrolyzer relative to the local renewable capacity and surrounding transmission network is of great importance, as a large electrolyzer capacity might induce congestion in the system, amplifying the negative effects on the cost of generation and the renewable power curtailment. Modeling the full transmission network when evaluating the presence of an electrolyzer in the power system might therefore be important because the effects of a congested network are otherwise unknown. However, if the REP is price sensitive both in theory (bidding its marginal cost curve) and in practice (marginal cost proportional to the system), these negative effects are largely mitigated.
This work leaves several future research directions. The REP might be exposed to uncertainty in its renewable production as well as to hydrogen demand constraints, which will affect the opportunity cost of selling electricity. Further, the REP’s potential as a price maker, and to what extent it might act strategically rather than bidding its true opportunity cost curve, can be analyzed. Finally, this work can be extended to investigate the impact of renewable hydrogen regulations and subsidies on system and REP results.